NEXT
Make New Entry,
Make Followup Entry
User name Jin Huang
Log entry time 00:09:07 on November 03, 2008
Entry number 246338
keyword=Target Expert Summary for Nov 02 Day-Swing Shift
Time Line & Link to Reports:
- 10:29 : Pumping Laser OFF
- 10:31 :
Optics Position, Vertical Holding Field, Long. RF NMR (by
Yi Zhang) .
Pol=~60%
- 10:39 :
Pick up Position, Vertical Holding Field, Long. RF NMR (by
Yi Zhang). The NMR signal
on pumping chamber is confirmed not changing much.
- 10:49 :
Pick up Position, Long Holding Field, Vert. RF NMR (by Yi
Zhang). It shows that Upstream
= 240mV; Downstream = 250 mV. => Pol = 0.12%/mV * (UpStream
NMR+ DownStream NMR)
- Unknown Time : rotate holding field to Longitudinal+
- 12:02 :
Pick up Position, Long Holding Field, Vert. RF NMR (by Yi
Zhang)
: Pol=53%.
- 12:04 :
Requested 5uA on 3He target with 4x4 raster for production
running.
- 13:44 :
An NMR measurement without lowering to pickup coil position
(by Yi Zhang) :
No reliable NMR signal is
collected, because No
pickup coil is adjusted to work with beam line level long.
polarization.
- 14:50 : An
NMR measurement without lowering to pickup coil position
(by Yi Zhang) : No reliable
NMR signal is collected.
- 15:24 & 15:30 :
two NMR measurement without lowering to pickup coil
position (by Yi Zhang) :
No reliable NMR signal is
collected.
- 15:58 :
Pick up Position, Long Holding Field, Vert. RF NMR
(by Jin Huang and Yi Zhang) :
Pol = 38%
- 18:08 :
Pick up Position, Long Holding Field, Vert. RF NMR (by Jin
Huang) :
Pol = 31%
- 19:00 :
Pick up Position, Long Holding Field, Vert. RF NMR (by Jin
Huang) :
Pol = 28%
- 19:10 : Start Moller Measurement.
- 19:20 : Couple test, Vertical
Laser On
- 19:20 :
Rotate to Vertical
+.
- 19:26 :
Empty Location, Vertical Holding Field, Long RF NMR
(by Jin Huang) : Pol = 24%
- 19:35 : EPR Measuremant get Pol =
25.9%
- 19:37 :
Empty Location, Vertical Holding Field, Long RF NMR
(by Jin Huang) : Pol = 25%
- 19:40 :
Keep Vertical Pumping
- 20:00 :
Change Vertical Fiber -> Duo Laser; Trans Fiber ->
Fap Laser
- 21:20 : Start Transverse Pumping
- 21:30 :
We get EPR FM sweep signal with only 30mV on diode. Among 3
major resonance,
Potassium resonance is @ 16.0MHz
- 21:42 :
Get EPR Pol = 28.2%. => Pol. = 2.16 * Vertical NMR
- 21:59 :
NMR Pol = 30%
- 22:21 :
NMR Pol = 33.5%
- 23:02 :
NMR Pol = 38 %
- 23:04 : Laser Off; Rotate -> Longitudinal+
holding Field; Target -> Pickup Location
- 23:10 -> 23:13 : 3 Junk NMR by Jin Huang:
Using Longi RF & No Field Sweep (sweep start stop
field setting is same);
No Field Sweep;
Wrong Software Backgroud Substraction
- 23:14 :
Pick up Position, Long Holding Field, Vert. RF NMR (by Jin
Huang) :
Pol = 31% (drop a lot!)
- 23:22 :
Target -> BeO Position
- 23:15 : Since polarization is low and Xin needs
additional H2 run. Target
is restored to transverse position to keep
pumping.
- 23:31 :
NMR Beo Position : Long RF : Trans Holding Field: Vert pick
up coil (by Jin Huang):
Pol
= 35% (presuming cali constant is same for
BeO & Empty Target)
- 23:58 :
NMR Beo Position : Long RF : Trans Holding Field: Vert pick
up coil (by Jin Huang):
Pol
= 38% (presuming cali constant is same for
BeO & Empty Target)
- 24:00 : Keep Trans Pumping
Calib Constants:
Vertical Holding Field, Long. RF, Pumping Chamber Trans
Coil:
Reference to HaLog 246037 11/01/08 22:48
Yi Qiang EPR calibration of vertical
pumping
Pol = 1.5%/mV * Trans NMR
Long Holding Field, Vert RF, Target Chamber Trans pickup
Coil:
Reference to HaLog 246271 11/02/08 16:29
jinhuang Before taking beam@4PM, he3
Pol. is estimated to be
38%.
Pol = 0.12%/mV * (UpStream NMR+
DownStream NMR)
Trans Holding Field, Long RF, Target Chamber Vert pickup
Coil:
Reference to HaLog
http://www.jlab.org/~adaq/halog/html/0811_archive/081102221509.html
.
Pol = 2.16%/mV * Vert NMR
A copy of this log entry has been emailed to: jpchen@jlab.org , yqiang@jlab.org, brads@jlab.org , averett@jlab.org, jinhuang@jlab.org